OpenStudy (anonymous):

is sin(x+y) +sin(x-y)=2cosxsiny an identity?

OpenStudy (jdoe0001):

add them up :)

OpenStudy (jdoe0001):

$$ \sin(\color{red}x)\cos(\color{blue}y) + \cos(\color{red}x)\sin(\color{blue}y) + \sin(\color{red}x)\cos(\color{blue}y) - \cos(\color{red}x)\sin(\color{blue}y)\\ \implies \color{green}{\pmatrix{\sin(\color{red}x)\cos(\color{blue}y) + \sin(\color{red}x)\cos(\color{blue}y)}} +\color{green}{\pmatrix{\cos(\color{red}x)\sin(\color{blue}y)-\cos(\color{red}x)\sin(\color{blue}y)}} $$ what do you think?

OpenStudy (anonymous):

i don't understand what you mean?